# Domain-Specific Translation with Open-Source Large Language Models: Resource-Oriented Analysis

### v1

@@ A.2.1 Mistral
A.2.2 NLLB
Similarly, when fine-tuning NLLB-200 3.3, we tried a few learning rates. Unlike Mistral, the learning rate 1e-3 did not achieve the best result. However, the learning rate 5e-5 revealed better evaluation scores, as illustrated by Table 12. Hence, we used it when fine-tuning NLLB-200 3.3 for all the language pairs.
@@ A.2.2 NLLB
A.3 Dataset Packing
In all the experiments in the main paper, we use packing=True for the SFTTrainer. Hence, we set add_eos_token to False as the trainer’s dataset packing is True. By default, SFTTrainer adds eos_token if dataset packing=True, so there is no need to add it manually. If packing is False, then add_eos_token should be set to True. The evaluation results use CTranslate2 with int8 quantization and are summarized in Table 13.
@@ A.3 Dataset Packing
A.4 CTranslate2 vs vLLM
@@
Lang Domain Context Library BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical zero-shot CTranslate2 30.54 54.76 62.27 48.89
@@
EN-PT medical zero-shot CTranslate2 33.53 60.09 55.39 63.22
@@
EN-SW medical zero-shot CTranslate2 0.67 10.88 312.19 25.89
@@
generic zero-shot CTranslate2 1.72 12.74 289.67 29.64
@@
SW-EN medical zero-shot CTranslate2 11.84 33.31 98.46 60.11
@@
generic zero-shot CTranslate2 16.77 36.99 95.29 60.03
@@
Table 14: Evaluation of Mistral 7B performance (baseline model without quantization) using CTranslate2 and vLLM inference engines, showing comparable results.
A.5 Quantization
We compared three cases of quantization when working with NLLB-200 CTranslate2, i.e. int8, float16, and no quantization (cf. Table 15). Models with int8 quantization have the smallest size, and they excel on CPU setups, while models with float16 quantization are smaller than non-quantized models, but larger than int8 models. Float16 models are faster on GPU setups. When it comes to translation performance, the differences between the three setups are marginal. All NLLB-200 3.3 results in the main paper are for non-quantized models.
Moreover, we compared generation with Llama-3 70B with and without Activation-aware Weight Quantization (AWQ) using vLLM. As Table 16 shows, results without quantization are slightly better. All the results in the main paper for LLMs are for non-quantized models, except Llama-3.1 405B, due to commutating constrains. Hence, we tested Llama-3 70B with and without quantization and reported the results in this section to understand the effect of AWQ on the translation quality.
Lang Domain Context Quantization BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical zero-shot None 41.37 64.10 52.11 59.32
Float16 41.30 64.08 52.15 59.22
Int8 41.33 64.07 52.17 59.09
one-shot None 43.04 65.55 54.49 55.84
Float16 43.00 65.54 54.62 55.83
Int8 43.03 65.31 54.41 55.44
EN-PT medical zero-shot None 39.32 64.98 48.19 81.77
Float16 39.30 64.97 48.21 81.81
Int8 39.13 64.89 48.35 81.87
one-shot None 39.79 64.36 48.38 79.04
Float16 39.79 64.36 48.38 79.05
Int8 39.74 64.33 48.42 78.95
EN-SW medical zero-shot None 32.61 58.24 53.29 82.85
Float16 32.87 58.41 53.08 82.86
Int8 32.88 58.36 53.08 83.03
one-shot None 34.06 59.12 51.57 82.37
Float16 34.07 59.23 51.45 82.38
Int8 33.97 59.15 51.41 82.42
generic zero-shot None 35.57 57.55 54.91 83.90
Float16 35.54 57.56 54.93 83.92
Int8 35.48 57.46 55.00 83.85
one-shot None 33.79 55.93 56.34 83.23
Float16 33.86 55.96 56.32 83.24
Int8 33.73 55.90 56.22 83.27
SW-EN medical zero-shot None 44.58 64.73 42.71 84.11
Float16 44.48 64.69 42.71 84.17
Int8 44.28 64.60 43.03 84.11
one-shot None 43.66 64.03 42.85 83.92
Float16 43.66 64.01 42.86 83.89
Int8 43.35 63.72 43.23 83.78
generic zero-shot None 42.87 61.59 48.69 82.00
Float16 42.89 61.56 48.63 82.01
Int8 42.83 61.53 48.59 82.00
one-shot None 41.38 59.99 50.85 80.92
Float16 41.40 60.01 50.89 80.95
Int8 41.35 60.01 50.72 80.97
Table 15: Comparison of translation performance of NLLB-200 3.3, across different quantization methods, using CTranslate2. The results show comparable performance.
Lang Context Quantization BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑

### v2

@@ B.2.1 Mistral
B.2.2 NLLB
Similarly, when fine-tuning NLLB-200 3.3B, we tried a few learning rates. Unlike Mistral, the learning rate 1e-3 did not achieve the best result. However, the learning rate 5e-5 revealed better evaluation scores, as illustrated by Table 11. Hence, we used it when fine-tuning NLLB-200 3.3B for all the language pairs.
B.3 Dataset Packing
In all the experiments in the main paper, we use packing=True for the SFTTrainer. Hence, we set add_eos_token to False as the trainer’s dataset packing is True. By default, SFTTrainer adds eos_token if dataset packing=True, so there is no need to add it manually. If packing is False, then add_eos_token should be set to True. The evaluation results are presented in Table 12.
@@ B.3 Dataset Packing
Appendix C Appendix: Inference
C.1 CTranslate2 vs vLLM
@@
Lang Domain Context Engine BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical zero-shot CT2 30.54 54.76 62.27 48.89
@@
EN-PT medical zero-shot CT2 33.53 60.09 55.39 63.22
@@
EN-SW medical zero-shot CT2 0.67 10.88 312.19 25.89
@@
generic zero-shot CT2 1.72 12.74 289.67 29.64
@@
SW-EN medical zero-shot CT2 11.84 33.31 98.46 60.11
@@
generic zero-shot CT2 16.77 36.99 95.29 60.03
@@
Table 13: Evaluation of Mistral 7B performance (baseline model without quantization) using CTranslate2 (CT2) and vLLM inference engines, showing comparable results.
C.2 Quantization
We compared quantization formats when working with NLLB-200 CTranslate2, i.e. int8, float16, no quantization (cf. Table 15). Models with int8 quantization have the smallest size, and they excel on CPU setups, while models with float16 quantization are smaller than non-quantized models, but larger than int8 models. Float16 models are faster on GPU setups. When it comes to translation performance, the differences between the three setups are marginal. All NLLB-200 3.3B results in the main paper are for non-quantized models.
Moreover, we compared generation with Llama-3 70B with and without Activation-aware Weight Quantization (AWQ) using vLLM. As Table 14 shows, results without quantization are slightly better. All the results in the main paper for LLMs are for non-quantized models, except Llama-3.1 405B and DeepSeek V3 685B, due to commutating constrains. Hence, we tested Llama-3 70B with and without quantization and reported the results in this section to understand the effect of AWQ on the translation quality.
Lang Context Quant. BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
@@ C.2 Quantization
NLLB-200 3.3B Mistral 7B
Lang Domain Context Quant. BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑ BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR medical zero-shot None 41.37 64.10 52.11 59.32 30.54 54.76 62.27 48.89
Float16 41.30 64.08 52.15 59.22 30.47 54.75 62.28 48.51
Int8 41.33 64.07 52.17 59.09 30.03 54.18 63.85 47.00
one-shot None 43.04 65.55 54.49 55.84 46.18 66.18 47.59 69.25
Float16 43.00 65.54 54.62 55.83 46.15 66.16 47.90 69.27
Int8 43.03 65.31 54.41 55.44 44.71 65.84 48.33 69.13
EN-PT medical zero-shot None 39.32 64.98 48.19 81.77 33.53 60.09 55.39 63.22
Float16 39.30 64.97 48.21 81.81 33.52 60.06 55.50 63.23
Int8 39.13 64.89 48.35 81.87 32.62 59.78 56.59 59.77
one-shot None 39.79 64.36 48.38 79.04 37.63 62.92 49.26 77.99
Float16 39.79 64.36 48.38 79.05 37.63 62.91 49.34 77.98
Int8 39.74 64.33 48.42 78.95 37.49 62.82 49.33 78.25
EN-SW medical zero-shot None 32.61 58.24 53.29 82.85 0.67 10.88 312.19 25.89
Float16 32.87 58.41 53.08 82.86 0.65 10.94 315.29 26.10
Int8 32.88 58.36 53.08 83.03 0.49 10.22 290.16 25.34
one-shot None 34.06 59.12 51.57 82.37 1.09 13.85 326.60 22.35
Float16 34.07 59.23 51.45 82.38 1.08 14.11 334.16 22.30
Int8 33.97 59.15 51.41 82.42 0.97 13.18 283.04 21.87
generic zero-shot None 35.57 57.55 54.91 83.90 1.72 12.74 303.84 29.64
Float16 35.54 57.56 54.93 83.92 1.72 12.83 309.21 29.68
Int8 35.48 57.46 55.00 83.85 1.30 11.50 278.97 28.16
one-shot None 33.79 55.93 56.34 83.23 2.49 15.96 338.57 27.19
Float16 33.86 55.96 56.32 83.24 2.50 16.23 344.52 27.35
Int8 33.73 55.90 56.22 83.27 1.99 14.45 294.13 25.65
SW-EN medical zero-shot None 44.58 64.73 42.71 84.11 12.25 33.46 93.90 60.06
Float16 44.48 64.69 42.71 84.17 12.19 33.37 95.38 60.18
Int8 44.28 64.60 43.03 84.11 11.15 32.79 102.77 59.75
one-shot None 43.66 64.03 42.85 83.92 16.35 37.66 85.37 66.24
Float16 43.66 64.01 42.86 83.89 16.46 37.67 84.67 66.29
Int8 43.35 63.72 43.23 83.78 16.99 37.72 82.32 66.10
generic zero-shot None 42.87 61.59 48.69 82.00 16.78 36.97 95.5 59.99
Float16 42.89 61.56 48.63 82.01 16.76 36.92 95.21 60.06
Int8 42.83 61.53 48.59 82.00 17.25 36.82 92.17 59.62
one-shot None 41.38 59.99 50.85 80.92 21.82 40.63 79.79 63.55
Float16 41.40 60.01 50.89 80.95 22.03 40.75 79.05 63.54
Int8 41.35 60.01 50.72 80.97 21.50 40.17 80.12 63.12
Table 15: Evaluation of the translation performance of NLLB-200 3.3B and Mistral 7B, both without quantization and with lower-precision quantization formats (float16 and int8) using CTranslate2. The results indicate similar performance across all formats.
Appendix D Appendix: Other Models
D.1 Llama 3.3 vs. Llama 3
We experimented with the Llama-3.3 70B instruction-tuned model, as it was promoted to be more efficient with “enhanced” quality (cf. the official model card). However, our evaluation does not show any considerable improvement compared to Llama-3 70B (cf. Table 16), probably due to the different types of tasks they target.
Lang Model Context BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR Llama-3 70B zero-shot 39.74 62.51 53.36 66.75
one-shot 51.93 70.57 43.03 76.76
Llama-3.3 70B zero-shot 39.73 62.99 55.22 69.16
one-shot 48.72 69.53 48.17 76.98
EN-PT Llama-3 70B zero-shot 43.86 67.35 43.95 84.05
one-shot 45.95 68.82 41.68 86.58
Llama-3.3 70B zero-shot 38.24 64.56 54.99 78.2
one-shot 45.25 68.42 42.6 86.28
Table 16: Comparison of translation performance of Llama-3 70B and Llama-3.3 70B instruct. In general, Llama-3 70B shows better performance.